A manufactured product has the following information for June.

Standard Actual
Direct materials (6 lbs. @ $8 per lb.) 50,400 lbs. @ $8.20 per lb.
Direct labor (3 hrs. @ $16 per hr.) 24,700 hrs. @ $16.50 per hr.
Overhead (3 hrs. @ $11 per hr.) $ 279,800
Units manufactured 8,300

1) Compute the standard cost per unit and the total cost variance for June. Indicate whether the cost variance is favorable or unfavorable.

Direct materials
Direct labor
Overhead
Total $0
Total cost variance

2)Compute the direct materials price variance and the direct materials quantity variance. Indicate whether each variance is favorable or unfavorable.

Actual Cost Standard Cost
$0 0 $0
$0
0

3)

Compute the direct labor rate variance and the direct labor efficiency variance. Indicate whether each variance is favorable or unfavorable.

Actual Cost 0 0 Standard Cost
$0 0 $0
$0
0
Answer #1

Solution 1:

Standard cost per unit
Direct materials $48.00
Direct labor $48.00
Overhead $33.00
Total $129.00

Total cost variance = Total standard cost - Total actual costs = (8300*$129) - ($413,280 + $407,550 + $279,800)

= $1,070,700 - $1,100,630 = $29,930 U

Solution 2:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
50400 $8.20 $413,280.00 50400 $8.00 $403,200.00 49800 $8.00 $398,400.00
$10,080.00 U $4,800.00 U
Direct Material Price Variance Direct Material Qty variance

Solution 3:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
24700 $16.50 $407,550.00 24700 $16.00 $395,200.00 24900 $16.00 $398,400.00
$12,350.00 U $3,200.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance